COULD SOMEONE PLEASE HELP ME ASAP? Write the equation of the line passing through the given point and having the given slope, using the point slope form y-y1=m(x-x1).write your answer in standard form ax=by=c,a>0.
HUH? how do I do this? I am really confused on this.
(-1,5);m=-4/7

Found 2 solutions by mananth, Dana Hamdan:
Answer by mananth(16946)

(-1,5);m=-4/7
y = mx b
plug x,y & m
5= (-4/7)*-1)+b
5=4/7 +b
5-4/7 =b
31/7 =b
..
y= -4/7 *x +31/7
*7
7y=-4x+31
..
4x+7y-31=0
..
m.ananth@hotmail.ca

Answer by Dana Hamdan(2)

this is a real easy question it's giving you a point in which the line passes through, and it's giving you it's slope. So all you have to do is plug in your info in the ewuation you are given.
so plug in
y-5=-4/7(x-(-1))
y-5=-4/7(x+1)
y-5=4/7x-4/7
+5.......+5
(((((y= 4/7x+31/7)))) then change it into standered form so move the 4/7x with the (y) so both varubales are in the same side.
<<<<<<>>>>> this is your answer :D